​​10Consider a buffer prepared by mixing 15 mL of a .300 M lactic acid (Ka= 1.4x10^-4) and 15 mL of .300 M potassium lactate. Calculate the pH.

2) Calculate the pH of a buffer when 15mL of a 0.150 M sodium hydroxide is mixed with 15 mL of a .300 M lactic acid
Answer #1

pH = pKa + log [potassium lactate] / [lactic acid]

pKa = - log Ka = - log [1.4 x 10-4] = 3.85

[lactic acid] = (15 x 0.3 / 30) = 0.15 M

[potassium lactate] = (15 x 0.3 / 30) = 0.15 M

pH = 3.85 + log [0.15]/[0.15]

pH = 3.85

2) millimoles of lactic acid = 15 x 0.3 = 4.5

millimoles of NaOH = 15 x 0.15 = 2.25

4.5 - 2.25 = 2.25 millimoles lactic acid left

2.25 millimoles sodium lactate formed

[sodium lactate] = 2.25 / 30 = 0.075 M

[lactic acid] = 2.25 / 30 = 0.075 M

pH = pKa + log [sodium lactate] / [lactic acid]

pH = 3.85 + log [0.075]/[0.075]

pH = 3.85